OpenStudy (anonymous):

red and grey bricks were used to build a decorative wall. the number of redbricks over the number of grey bricks was 5/2. there were 175 bricks used in all. how many red bricks were used?

OpenStudy (anonymous):

could someone explain this to me like step by step

OpenStudy (anonymous):

I am in desperate need of help!

OpenStudy (lucaz):

@hartnn help?

hartnn (hartnn):

"the number of redbricks over the number of grey bricks was 5/2" so let the number of red bricks be 5x and grey bricks be 2x total = 175 so, 5x+2x =175 can you find x from here, ??

OpenStudy (anonymous):

i got x from that which is 25 but isnt there more to solving to see how many red bricks @hartnn

hartnn (hartnn):

exactly x= 25 is correct :) number of red brick = 5x = ..

OpenStudy (anonymous):

then to get total number of red bricks wouldnt you do 25x5 to get total

hartnn (hartnn):

yes, total red bricks = 5*25 = 125 :)

OpenStudy (anonymous):

thanks

hartnn (hartnn):

welcome ^_^